OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (anonymous):

answer is x=2

OpenStudy (anonymous):

-3(x+3) = -3(x+1)-5 Distribute the -3's on both sides. -3x-9 = -3x -3 - 5 add 3x to both sides. add 9 to both sides and simplify. 0x=1 Since you can't divide by a 0; there is no solution. does that help?